OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is 5x-2y = 11?

OpenStudy (anonymous):

5/2 y=(5/2)x-(11/2)

OpenStudy (anonymous):

just the slope? m = 5/2

OpenStudy (anonymous):

this is easy basic algebra. do some reading or online research to figure this out first then come here. dont clutter our feed with your whole homework assignment kid. TRUST ME start leanring now, youll think yourself later!